OpenStudy (anonymous):

Identify the slope in the following equation y = 2x - 12. -2 2 12 -12

OpenStudy (anonymous):

m=2

jimthompson5910 (jim_thompson5910):

y = 2x-12 is in the form y = mx+b where m is the slope. We can see that m = 2. So the slope is 2.

OpenStudy (anonymous):

thank you guys!! ;D

OpenStudy (anonymous):

jules <3 Best Answer - Chosen by Asker y=2x+ -12 slope is 2 y intercept is 12 slope intercept form is y=mx+b m is slope b is y intercept Source(s): my brain